OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = quantity x minus nine divided by quantity x plus five. and g(x) = quantity negative seven x minus eight divided by quantity x minus one.

OpenStudy (anonymous):

\[f \left( x \right)=\frac{ x-9 }{ x+5} \text{ , }g \left( x \right)=\frac{ -7x-8 }{ x-1 }\]like this?

OpenStudy (anonymous):

do you know how to do f(g(x)) and g(f(x))?